• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Nathan

    Log entry time 00:14:39 on January13,2002

    Entry number 80160

    keyword=Shift summary

    At the start of the shift, the target group said they were finishing up their work on the slow controls for the cryogenic target. By 2000, they left the target in the hands of J.P. Chen, who ran into lots of problems (see his report in a later log). As the shift ended, the hall is still in Controlled Access and J.P. and Dave Meekins are still working on the target. No beam into the hall during this shift.